OpenStudy (anonymous):

Please help me with this problem: 1.6q-5.7-2.9q=-2.3q-3.8.

OpenStudy (anonymous):

The first step is to add like terms on each side of the = sign. The only thing you can add is the 1.6q - 2.9q and you would get -1.3q So you would have -1.3q - 5.7 = -2.3q = 3.8 Do you understand this so far?

OpenStudy (anonymous):

I'm trying to work the problem with you and I got -1.3q-5.7=-2.3q+3.8

OpenStudy (anonymous):

Good.. now I always like to get rid of the smallest "x" if I have one on both sides. So let's get rid of the -2.3q. Try it and let me know what you get.

OpenStudy (anonymous):

q

OpenStudy (anonymous):

Yes you get q - 5.7 = 3.8 Then finish and let me know what you get.

OpenStudy (anonymous):

q=9.5

OpenStudy (anonymous):

Is q=9.5 right?